# A central limit theorem for partitions involving generalised divisor function

### v1

@@ 2. The Main Results
where the constants $C_{\mu}(r)$ and $C_{\sigma}(r)$ are defined in (5.3) and (5.1), respectively, and $\eta$ is the implicit solutions of
@@ 2.1. Initial Setup
Since $\mu_{f,n}/\nu_{f,n}$ a is constant, we can write
@@
$\mathcal{M}\left[\sum_{k}\lambda_{k}h(\mu_{k}x);s\right]=h^{*}(s)\sum_{k}\frac{\lambda_{k}}{\mu_{k}}.$
@@ 3. The Dirichlet series
$c_{\chi}(n)\coloneqq\sum_{b(\bmod{m})}\chi(b)e\left(\frac{bn}{m}\right),\quad\text{and}\quad c^{\prime}_{\chi}(n)\coloneqq\sum_{\begin{subarray}{c}b(\bmod{m})\\ (b,m)=1\end{subarray}}\chi(b)e\left(\frac{bn}{m}\right).$ (3.5)
@@
where $c_{m}(n)$ denotes the Ramanujan sum. The following identity then holds for $\mathrm{Re}(s)>1$ or $\mathrm{Re}(s)+r>0$:
@@
$K_{r}(s)=C(r)\prod_{p}\left(1-A_{p,r}(s)\right)$ (3.9)
@@
$A_{p,r}(s)=\frac{\frac{1}{p^{r+1}(p-1)}}{1+\frac{1}{p^{r+1}(p-1)}}\left(\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\frac{1}{p^{2s+2r+2}}+\cdots-\frac{1}{p^{(k+1)s+r+1}}+\cdots\right)$
@@
$=$ $-\left(\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\cdots-\frac{1}{p^{(k+1)s+r+1}}+\cdots\right).$ (3.13)
@@
$|K_{r}(s)|\leq C(r)\prod_{p}\left(1-|A_{p,r}(s)|\right)\leq C(r)\prod_{p}\left(1-\frac{2^{\sigma_{0}+r+1}}{(2^{\sigma_{0}+r+1}-1)p^{\sigma_{0}+r+1}}\right).$
@@
$c^{\prime}_{\chi}(n)=\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}\chi(b)e\left(\frac{bn}{m}\right)=\sum_{\begin{subarray}{c}b\bmod{m}\end{subarray}}\chi_{0}(b)\chi(b)e\left(\frac{bn}{m}\right)=\sum_{\begin{subarray}{c}b\bmod{m}\end{subarray}}\chi_{1}(b)e\left(\frac{bn}{m}\right)=c_{\chi_{1}}(n),$ (3.14)
@@
$E_{r}(\sigma)=C^{\prime}(r)\prod_{p}(1-J_{p,r}(\sigma))$ (3.17)
@@
$J_{p,r}(\sigma)=\frac{\frac{1-p^{-r}}{p^{r+1}}\left(1-\frac{1}{2p^{3\sigma+2r+1}(1-2^{-\sigma+r+1})}\right)}{1+\frac{1-p^{-r}}{p^{r+1}}}$
@@
Let $\chi$ be a Dirichlet character modulo $m$, induced by a primitive character $\chi^{*}$ modulo $m^{*}$, where $m^{*}|m$. From (3.14), it is evident that $c^{\prime}_{\bar{\chi}}(n)=c_{\bar{\chi}_{1}}(n)$ with $\bar{\chi}_{1}=\chi_{0}\bar{\chi}=\chi_{0}\chi^{*}_{0}\bar{\chi}^{*}=\chi_{0}\bar{\chi}^{*}$.
@@
$c^{\prime}_{\bar{\chi}}(n)c_{{\chi}}(1)=$ $\chi^{*}\left(\frac{n}{(m,n)}\right)\bar{\chi}^{*}\left(\frac{m/(m,n)}{m^{*}}\right)\mu\left(\frac{m/(m,n)}{m^{*}}\right)\frac{\varphi(m)}{\varphi(m/(m,n))}\mu\left(\frac{m}{m^{*}}\right)\chi^{*}\left(\frac{m}{m^{*}}\right)c_{\bar{\chi}^{*}}(1)c_{{\chi^{*}}}(1)$
@@
$\frac{c^{\prime}_{\bar{\chi}}(n)c_{{\chi}}(1)}{\varphi(m)}=$ $m^{*}\chi^{*}\left(\ell\right)\bar{\chi}^{*}\left(k\right)\mu\left(k\right)\frac{1}{\varphi(m^{*}k)}\mu\left(\delta k\right)\chi^{*}\left(\delta k\right)$
@@
$D_{2}(s,r)=$ $\sum_{m=1}^{\infty}\frac{1}{m^{r+1}}\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\sum_{n=1}^{\infty}\frac{c_{\chi}(1)c^{\prime}_{\bar{\chi}}(n)}{\varphi(m)n^{s}}$
@@
$\frac{L(s+r+1,\chi^{*})}{L(s+r,\chi^{*})}\prod_{\begin{subarray}{c}p|k\\ p\nmid m^{*}\end{subarray}}(1-B_{p,r}(s,\chi^{*}))\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right)^{-1},$
@@
$B_{p,r}(s,\chi^{*})=\frac{\left(\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r)+1}}-\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r+1)}}+\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r)+1}}-\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r+1)}}+\cdots\right)}{\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}-\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r)+1}}+\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r+1)}}-\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r)+1}}+\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r+1)}}-\cdots\right)}.$ (3.22)
@@
$A_{r}(s,\chi^{*})=\prod_{p|k}(1-B_{p,r}(s,\chi^{*}))\left(1-\frac{\chi^{*}(p)}{p^{s}}\right)\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right)^{-1}.$ (3.23)
@@
$\sum_{p|k}|1-B_{p,r}(s,\chi^{*})|<\frac{1}{2}\sum_{p|k}\frac{1}{p^{2\sigma+2r+1}}<\infty.$
@@
$\left|\frac{\chi^{*}(p)}{p^{s}}\right|\leq\left|\frac{1}{p^{s}}\right|=\frac{1}{p^{\sigma}}\quad\text{and}\quad$ $\left|1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right|\geq 1-\frac{1}{2^{\sigma+r+1}}.$
@@
$|A_{r}(s,\chi^{*})|\leq$ $\prod_{p|k}|1-B_{p,r}(s,\chi^{*})|\left|\frac{1}{p^{s}}\right|\left|1+\frac{1}{p^{s+r+1}}-\frac{1}{p^{s+r}}\right|^{-1}\leq\prod_{p|k}\frac{1}{2p^{3\sigma+2r+1}\left(1-\frac{1}{2^{\sigma+r+1}}\right)}.$ (3.25)
Therefore, $|A_{r}(s,\chi^{*})|$ is uniformly bounded in the half-plane $\mathrm{Re}(s)>-2r/3$. Moreover, since
@@
$\leq$ $\varphi(m^{*})\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)}{\zeta(2(\sigma+r))}\prod_{p|k}\frac{1}{2p^{3\sigma+2r+1}\left(1-\frac{1}{2^{\sigma+r+1}}\right)}.$
Inserting this result, we obtain
$|D_{2}(s,r)|\leq\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)}{\zeta(2(\sigma+r))}\sum_{k\geq 1}\frac{\mu^{2}(k)}{k^{r+1}}\prod_{p|k}\frac{1}{2p^{3\sigma+2r+1}\left(1-\frac{1}{2^{\sigma+r+1}}\right)}\sum_{\begin{subarray}{c}m^{*}\geq 1\\ (m^{*},k)=1\end{subarray}}\frac{1}{{m^{*}}^{r}}.$
@@
$|D_{2}(s,r)|\leq\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)\zeta(r)}{\zeta(2(\sigma+r))}\prod_{p}\left(1+\frac{1-p^{-r}}{p^{r+1}}\frac{1}{2p^{3\sigma+2r+1}(1-\frac{1}{2^{\sigma+r+1}})}\right),$
@@
$E_{r}(\sigma)=\prod_{p}\left(1+\frac{1-p^{-r}}{p^{r+1}}\frac{1}{2p^{3\sigma+2r+1}(1-\frac{1}{2^{\sigma+r+1}})}\right).$
@@
$E_{r}(\sigma)=C^{\prime}(r)\prod_{p}\frac{\left(1+\frac{1-p^{-r}}{p^{r+1}}(1-E_{p,r}(\sigma))\right)}{1+\frac{1-p^{-r}}{p^{r+1}}},$
@@
$E_{p,r}(\sigma)=1-\frac{1}{2p^{3\sigma+2r+1}(1-\frac{1}{2^{\sigma+r+1}})}.$
From our analysis of $B_{p,r}(s,\chi^{*})$, it follows that $E_{p,r}(\sigma)$ converges absolutely for $\sigma>-2r/3$, thereby completing the proof of the lemma. ∎
@@ 4.1. Estimate of $\Omega_{1,f}$
where $\mathrm{Re}(s)=\sigma$ and $E_{r}(\sigma)$ is defined in (3.17). The series is absolutely convergent for $\mathrm{Re}(s)-i-j+2r/3>0$. However, triple poles arise from the zeta functions at $\sigma=i+j+1$, $\sigma=i+j+1-r$, and $\sigma=i+j-r$. By invoking Theorem 2.3, we express this as
@@
11footnotetext: Note that $E_{r}(-r)$ and $E_{r}(1-r)$ are constants.
@@
$\left|-i\sum_{k\leq k_{0}}^{\infty}\frac{k^{3}\Delta_{f}(k)(1-u^{-1}e^{k(\tau+i\theta)})}{(1+u^{-1}e^{k(\tau+i\theta)})^{3}}\right|\leq\sum_{k\leq k_{0}}\frac{k^{3}|\Delta_{f}(k)|(1+ve^{k\tau})}{|1+ve^{k(\tau+i\theta)}|^{3}}$
@@ 4.2. Estimate of $\Omega_{2,f}$
$\left(\frac{|Q_{\sigma_{k}}(e^{\tau+iy},u)|}{Q_{\sigma_{k}}(e^{-\tau,u})}\right)^{2}=$ $\prod_{n=1}^{\infty}\left(1-\frac{2ue^{-n\tau}(1-\cos(ny))}{(1+ue^{-n\tau})^{2}}\right)^{\Delta_{f}(n)}$
@@
$\left(\frac{|Q_{\sigma_{k}}(e^{\tau+iy},u)|}{Q_{\sigma_{k}}(e^{-\tau,u})}\right)^{2}\leq\exp\left(-\frac{2u(r-1)}{(1+u)^{2}}\varrho\left(\frac{e^{-\tau}}{(1-e^{-\tau})^{r}}-\frac{e^{-\tau}}{\left|1-e^{-\tau-iy}\right|^{r}}\right)\right).$ (4.15)
@@
$|\Omega_{2,\sigma_{r}}|\leq\frac{e^{n\tau}}{2\pi}\int_{\theta_{n}}^{\pi}\left|{Q}_{\sigma_{k}}(e^{-\tau-i\theta},u)\right|d\theta=\frac{e^{n\tau+F(\tau,u)}}{\pi}\int_{\theta_{n}}^{\pi}\frac{|Q_{\sigma_{k}}(e^{-\tau-i\theta},u)|}{Q_{\sigma_{k}}(e^{-\tau},u)}d\theta.$

### v2

@@ 2. The Main Results
where the constants $C_{\mu}(r)$ and $C_{\sigma}(r)$ are defined in (5.3) and (5.1), respectively, and $\eta$ is the implicit solution of
@@ 2.1. Initial Setup
Since $\mu_{f,n}/\nu_{f,n}$ is a constant, we can write
@@
$\mathcal{M}\left[\sum_{k}\lambda_{k}h(\mu_{k}x);s\right]=h^{*}(s)\sum_{k}\frac{\lambda_{k}}{\mu_{k}^{s}}.$
@@ 3. The Dirichlet series
$c^{\prime}_{\chi}(n)\coloneqq\sum_{\begin{subarray}{c}b(\bmod{m})\\ (b,m)=1\end{subarray}}\chi(b)e\left(\frac{bn}{m}\right).$ (3.5)
@@
where $c_{m}(n)$ denotes the Ramanujan sum. The following identity then holds for $\mathrm{Re}(s)>1$:
@@
$K_{r}(s)=C(r)\prod_{p}\left(1+A_{p,r}(s)\right)$ (3.9)
@@
$A_{p,r}(s)=-\frac{\frac{1}{p^{r+1}(p-1)}}{1+\frac{1}{p^{r+1}(p-1)}}\left(\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\frac{1}{p^{2s+2r+2}}+\cdots-\frac{1}{p^{(k+1)(s+r+1)}}+\cdots\right)$
@@
$=$ $-\left(\frac{1}{p^{s}}-\frac{1}{p^{s+r+1}}+\frac{1}{p^{2s+r+1}}-\cdots-\frac{1}{p^{(k+1)(s+r+1)}}+\cdots\right).$ (3.13)
@@
$|K_{r}(s)|\leq C(r)\prod_{p}\left(1+|A_{p,r}(s)|\right)\leq C(r)\prod_{p}\left(1+\frac{2^{\sigma_{0}+r+1}}{(2^{\sigma_{0}+r+1}-1)p^{\sigma_{0}+r+1}}\right).$
@@
$c^{\prime}_{\chi}(n)=\sum_{\begin{subarray}{c}b\bmod{m}\\ (b,m)=1\end{subarray}}\chi(b)e\left(\frac{bn}{m}\right)=\sum_{\begin{subarray}{c}b\bmod{m}\end{subarray}}\chi_{0}(b)\chi(b)e\left(\frac{bn}{m}\right)=\sum_{\begin{subarray}{c}b\bmod{m}\end{subarray}}\chi_{1}(b)e\left(\frac{bn}{m}\right)=c^{\prime}_{\chi_{1}}(n),$ (3.14)
@@
$E_{r}(\sigma)=C^{\prime}(r)\prod_{p}(1+J_{p,r}(\sigma))$ (3.17)
@@
$J_{p,r}(\sigma)=-\frac{\frac{1-p^{-r}}{p^{r+1}}\left(\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)}{1+\frac{1-p^{-r}}{p^{r+1}}}$
@@
Let $\chi$ be a Dirichlet character modulo $m$, induced by a primitive character $\chi^{*}$ modulo $m^{*}$, where $m^{*}|m$. From (3.14), it is evident that $c^{\prime}_{\bar{\chi}}(n)=c^{\prime}_{\bar{\chi}_{1}}(n)$ with $\bar{\chi}_{1}=\chi_{0}\bar{\chi}=\chi_{0}\chi^{*}_{0}\bar{\chi}^{*}=\chi_{0}\bar{\chi}^{*}$.
@@
$c^{\prime}_{\bar{\chi}}(n)c^{\prime}_{{\chi}}(1)=$ $\chi^{*}\left(\frac{n}{(m,n)}\right)\bar{\chi}^{*}\left(\frac{m/(m,n)}{m^{*}}\right)\mu\left(\frac{m/(m,n)}{m^{*}}\right)\frac{\varphi(m)}{\varphi(m/(m,n))}\mu\left(\frac{m}{m^{*}}\right)\chi^{*}\left(\frac{m}{m^{*}}\right)c^{\prime}_{\bar{\chi}^{*}}(1)c^{\prime}_{{\chi^{*}}}(1)$
@@
$\frac{c^{\prime}_{\bar{\chi}}(n)c^{\prime}_{{\chi}}(1)}{\varphi(m)}=$ $m^{*}\chi^{*}\left(\ell\right)\bar{\chi}^{*}\left(k\right)\mu\left(k\right)\frac{1}{\varphi(m^{*}k)}\mu\left(\delta k\right)\chi^{*}\left(\delta k\right)$
@@
$D_{2}(s,r)=$ $\sum_{m=1}^{\infty}\frac{1}{m^{r+1}}\sum_{\begin{subarray}{c}\chi(\bmod{m})\\ \chi\neq\chi_{0}\end{subarray}}\sum_{n=1}^{\infty}\frac{c^{\prime}_{\chi}(1)c^{\prime}_{\bar{\chi}}(n)}{\varphi(m)n^{s}}$
@@
$\frac{L(s+r+1,\chi^{*})}{L(s+r,\chi^{*})}\prod_{\begin{subarray}{c}p|k\\ p\nmid m^{*}\end{subarray}}(1+B_{p,r}(s,\chi^{*}))\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right)^{-1},$
@@
$B_{p,r}(s,\chi^{*})=-\frac{\left(\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r)+1}}-\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r+1)}}+\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r)+1}}-\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r+1)}}+\cdots\right)}{\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}-\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r)+1}}+\frac{{\chi^{*}}(p)^{2}}{p^{2(s+r+1)}}-\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r)+1}}+\frac{{\chi^{*}}(p)^{3}}{p^{3(s+r+1)}}-\cdots\right)}.$ (3.22)
@@
$A_{r}(s,\chi^{*})=\prod_{p|k}(1+B_{p,r}(s,\chi^{*}))\left(1-\frac{\chi^{*}(p)}{p^{s}}\right)\left(1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right)^{-1}.$ (3.23)
@@
$\sum_{p|k}|B_{p,r}(s,\chi^{*})|<\frac{1}{2}\sum_{p|k}\frac{1}{p^{2\sigma+2r+1}}<\infty.$
@@
$\left|1-\frac{\chi^{*}(p)}{p^{s}}\right|\leq\left|1+\frac{1}{p^{s}}\right|=1+\frac{1}{p^{\sigma}}\quad\text{and}\quad$ $\left|1+\frac{\chi^{*}(p)}{p^{s+r+1}}-\frac{\chi^{*}(p)}{p^{s+r}}\right|\geq 1-\frac{1}{2^{\sigma+r+1}}.$
@@
$|A_{r}(s,\chi^{*})|\leq$ $\prod_{p|k}|1+B_{p,r}(s,\chi^{*})|\left|1+\frac{1}{p^{s}}\right|\left|1+\frac{1}{p^{s+r+1}}-\frac{1}{p^{s+r}}\right|^{-1}$
$\leq$ $\prod_{p|k}\left(1+\frac{1}{2p^{2\sigma+2r+1}}\right)\left(1+\frac{1}{p^{\sigma}}\right)\left(1-\frac{1}{2^{(\sigma+r+1)}}\right)^{-1}$
$\leq$ $\left(1-\frac{1}{2^{(\sigma+r+1)}}\right)^{-1}\prod_{p|k}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}+\frac{1}{2p^{3\sigma+2r+1}}\right).$ (3.25)
Therefore, $|A_{r}(s,\chi^{*})|$ is uniformly bounded in the half-plane $\mathrm{Re}(s)>1$. Moreover, since
@@
$\leq$ $\varphi(m^{*})\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)}{\zeta(2(\sigma+r))}\prod_{p|k}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\left(1-\frac{1}{2^{\sigma+r+1}}\right)^{-1}.$
Ignoring the higher-order terms and inserting the expression above, we obtain
$|D_{2}(s,r)|\leq\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)}{\zeta(2(\sigma+r))}\sum_{k\geq 1}\frac{\mu^{2}(k)}{k^{r+1}}\prod_{p|k}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\sum_{\begin{subarray}{c}m^{*}\geq 1\\ (m^{*},k)=1\end{subarray}}\frac{1}{{m^{*}}^{r}}.$
@@
$|D_{2}(s,r)|\leq\frac{\zeta(\sigma)\zeta(\sigma+r)\zeta(\sigma+r+1)\zeta(r)}{\zeta(2(\sigma+r))}\prod_{p}\left(1+\frac{1-p^{-r}}{p^{r+1}}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\right),$
@@
$E_{r}(\sigma)=\prod_{p}\left(1+\frac{1-p^{-r}}{p^{r+1}}\left(1+\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right)\right).$
@@
$E_{r}(\sigma)=C^{\prime}(r)\prod_{p}\frac{\left(1+\frac{1-p^{-r}}{p^{r+1}}(1+E_{p,r}(\sigma))\right)}{1+\frac{1-p^{-r}}{p^{r+1}}},$
@@
$E_{p,r}(\sigma)=-\left(\frac{1}{p^{\sigma}}+\frac{1}{2p^{2\sigma+2r+1}}\right).$
From our analysis of $B_{p,r}(s,\chi^{*})$, it follows that $E_{p,r}(\sigma)$ converges absolutely for $\sigma>1$, thereby completing the proof of the lemma. ∎
@@ 4.1. Estimate of $\Omega_{1,f}$
where $\mathrm{Re}(s)=\sigma$ and $E_{r}(\sigma)$ is defined in (3.17). However, triple poles arise from the zeta functions at $\sigma=i+j+1$, $\sigma=i+j+1-r$, and $\sigma=i+j-r$. By invoking Theorem 2.3, we express this as
@@
11footnotetext: Note that due to the convergence region of $E_{r}(\sigma)$, $E_{r}(-r)$ and $E_{r}(1-r)$ are constants.
@@
$\left|-i\sum_{k\leq k_{0}}\frac{k^{3}\Delta_{f}(k)(1-u^{-1}e^{k(\tau+i\theta)})}{(1+u^{-1}e^{k(\tau+i\theta)})^{3}}\right|\leq\sum_{k\leq k_{0}}\frac{k^{3}|\Delta_{f}(k)|(1+ve^{k\tau})}{|1+ve^{k(\tau+i\theta)}|^{3}}$
@@ 4.2. Estimate of $\Omega_{2,f}$
$\left(\frac{|Q_{\sigma_{r}}(e^{\tau+iy},u)|}{Q_{\sigma_{r}}(e^{-\tau,u})}\right)^{2}=$ $\prod_{n=1}^{\infty}\left(1-\frac{2ue^{-n\tau}(1-\cos(ny))}{(1+ue^{-n\tau})^{2}}\right)^{\Delta_{f}(n)}$
@@
$\left(\frac{|Q_{\sigma_{r}}(e^{\tau+iy},u)|}{Q_{\sigma_{r}}(e^{-\tau,u})}\right)^{2}\leq\exp\left(-\frac{2u(r-1)}{(1+u)^{2}}\varrho\left(\frac{e^{-\tau}}{(1-e^{-\tau})^{r}}-\frac{e^{-\tau}}{\left|1-e^{-\tau-iy}\right|^{r}}\right)\right).$ (4.15)
@@
$|\Omega_{2,\sigma_{r}}|\leq\frac{e^{n\tau}}{2\pi}\int_{\theta_{n}}^{\pi}\left|{Q}_{\sigma_{r}}(e^{-\tau-i\theta},u)\right|d\theta=\frac{e^{n\tau+F(\tau,u)}}{\pi}\int_{\theta_{n}}^{\pi}\frac{|Q_{\sigma_{r}}(e^{-\tau-i\theta},u)|}{Q_{\sigma_{r}}(e^{-\tau},u)}d\theta.$